# HMD2: Environment-aware Motion Generation from Single Egocentric Head-Mounted Device

### v1

@@ Abstract
This paper investigates the online generation of realistic full-body human motion using a single head-mounted device with an outward-facing color camera and the ability to perform visual SLAM. Given the inherent ambiguity of this setup, we introduce a novel system, HMD2, designed to balance between motion reconstruction and generation. From a reconstruction standpoint, our system aims to maximally utilize the camera streams to produce both analytical and learned features, including head motion, SLAM point cloud, and image embeddings. On the generative front, HMD2 employs a multi-modal conditional motion Diffusion model, incorporating a time-series backbone to maintain temporal coherence in generated motions, and utilizes autoregressive in-painting to facilitate online motion inference with minimal latency (0.17 seconds). Collectively, we demonstrate that our system offers a highly effective and robust solution capable of scaling to an extensive dataset of over 200 hours collected in a wide range of complex indoor and outdoor environments using publicly available smart glasses.
@@ 1 Introduction
Wearable devices such as smart glasses promise to become the cornerstone of next-generation personal computing. A key challenge is accurately interpreting the wearer’s self-motion from the device’s limited input signals, taking into account the social and environmental context at the moment. The capability to generate full-body movements solely from a single head-mounted device (HMD) in real-time, outdoors and indoors, will open the door to many downstream applications, including telepresence, fitness and health monitoring, and navigation.
@@
In this paper, we introduce HMD2 (Head-conditioned Motion Diffusion from HMD), the first system, to our knowledge, capable of online generation of full-body movements from a single HMD (Project Aria Glasses [14]), conditioned on outward-facing egocentric camera streams in diverse environments. Such an input device can only provide limited observation of the body and surroundings, so the critical question is how to maximally utilize the input signals. Our insight is to reuse the input data to generate features of different modalities and to cover independent aspects of the environment and the motion maximally. Specifically, from the input streams, we mix and match analytical and learning toolboxes to extract 1. wearer’s head motion from off-the-shelf real-time visual SLAM; 2. environment feature points as a by-product of SLAM, important for motion disambiguation in complex scenes; and 3. head camera image embeddings (e.g. using CLIP [56]) for additional scene clues and intermittently visible body parts.
However, full recovery of the wearer’s motion is still highly under-constrained, given our input. Our system takes a generative approach and adopts a Diffusion-based Transformer backbone to strike a balance between motion reconstruction and motion generation. This allows for the generation of diverse outcomes, such as varying leg movements, from identical inputs. Additionally, our Diffusion model can predict motions with minimal future sensor information (0.17 seconds), facilitating online and real-time use cases.
Contrary to evaluations using large synthetic datasets or small-scale real-world datasets, we train and test our system on the extensive 200-hour real-world Nymeria dataset [45] recorded with publicly available head-mounted device, containing various indoor and outdoor activities performed by over 100 subjects with diverse body sizes and demographics. While most existing research on motion tracking is evaluated by only reconstruction accuracy, we acknowledge the inherent ambiguity in our problem and evaluate our system on generation fidelity and diversity as well. Our contributions are summarized as follows:
@@
    We demonstrate the adaption of a time-series motion Diffusion model for online autoregressive inference through in-painting, eliminating the dependency on future sensor input and achieving minimal latency.
@@ 2 Related Work
Human Motion from Sparse Sensors.  Capturing motion with wearable sensors has gained interest across fields like Computer Vision, Graphics, and Health. Self-contained sensors like IMUs [66], electromagnetic sensors [35], and EMGs [10] offer motion reconstruction without the need for costly studios with multiple cameras. The sparse placement of these sensors on the body, aimed at reducing user friction, coupled with high noise levels in most sensors, invites the use of learning methodologies to enhance reconstruction quality. Examples include configurations employing six IMUs [66, 28, 76, 33, 75], and head and wrists VR trackers [32, 13, 88, 7, 73, 31].
Our approach focuses on using a single wearable device to minimize user friction, though this increases the challenge of ”recovering” full motion. However, for many applications like telepresence, one would prioritize visually appealing, realistic, and diverse inferred motions rather than precise. Therefore, we evaluate our system not just on reconstruction accuracy but also on realism and diversity – metrics often overlooked in this field.
Pose and Motion from Egocentric Cameras.  Wearable egocentric cameras are a natural choice for self-contained motion generation systems, which saw increasing research interest and is closely related to our work. Two main types of body-mounted cameras have seen active research efforts - downward-facing (commonly fish-eye) cameras and outward-facing cameras. For downward cameras [57, 8, 61, 68, 72, 83, 85, 52, 34, 41], most works formulate the learning architecture as a pre-frame one, predicting current pose from corresponding camera image(s). While such design choice simplifies building an online system, they traded off temporal coherence in human motions. Wang et al. [67] tackled this issue by adopting a Diffusion model for temporal regularization in a separate refinement stage, which inspired us to adopt a Diffusion backbone and a single-stage time-window-based learning architecture. Both synthesized [4] and, more recently, real-device [3] datasets are used to train and evaluate such methods.
Outward-facing cameras are more common on current publicly available devices (e.g. Project Aria [14]), though egocentric motion generation is less explored in this setup. As pointed out by one of the seminal works using a chest-mounted camera [30], the main technical challenge is intermittent and partial visibility of the body, in contrast to camera distortion in the downward fish-eye setup. Later works [78, 79, 43] explored simulation-based techniques where physics helped fill in the missing information of motion, and Li et al. [39] showed for the first time such system’s generalizability to large-scale datasets. We build upon EgoEgo [39] and harness additional visual cues from the camera streams beyond inferring head poses, and additionally support non-flat terrains and low-latency long sequence generations.
@@
Learning-based Pose and Motion Generation.  Generating controllable and realistic human movements is a long-standing goal in Computer Graphics and Vision. Modern deep learning opens new possibilities for this problem, with earlier attempts exploring both regression-based [25, 24] and generative [21, 40] frameworks. Recently Diffusion models demonstrated impressive capabilities in the generative setting across various tasks such as text [58, 89], music [62], and audio [5] conditioned motion generation. While the field starts to see conditional Diffusion methods where the control signal is temporally dense [7, 13], frameworks that generate motions in an online fashion with minimal latency [64] are still under-explored. Our work adopts autoregressive inpainting for low-latency inference – the concept of autoregressive Diffusion models has been explored in the motion domain albeit in different contexts [19, 84, 59, 77].
The success of Diffusion models in motion synthesis has also intrigued researchers to use them for pose reconstruction, for example from third- and second-person images, especially when ambiguity exists [11, 12, 15, 16, 26, 82, 81]. Our task is similarly highly ambiguous, and our system adopts Transformer-based Diffusion models to help generate temporally coherent motions.
@@ 3 Method
Figure 1: Overview: HMD2 generates realistic full-body motion that aligns with the signals from a single head-mounted device. Using the image streams from the ego-centric camera and head trajectory with the feature cloud from the onboard SLAM system, we employ a diffusion-based framework to generate the wearer’s full-body motion.
@@
Given the under-constrained nature of the task, we employ a Diffusion model [22] with a time-series Transformer encoder [65] to model the motion distribution. To ensure temporal consistency during streaming, we use autoregressive in-painting during denoising, aligning new body motion with previous predictions.
@@ 3.1 Multi-modal Scene and Motion Conditions
Head Pose Trajectory.  The device pose provides precise spatial location and movement of the wearer’s head. For this pose condition, we augment the device pose vector with its linear and angular velocity vector $(\bm{v},\bm{\omega})$ computed from finite differences to form $\bm{p}=\{\bm{t},\bm{R},\bm{v},\bm{\omega}\}$. We canonicalize each window of $\{\bm{p}\}_{0,1,\cdots,T}$ to its first frame $\bm{p}_{0}$, which allows the model to function in arbitrarily large spaces and generate infinitely long sequences. This is crucial for navigation in a multistory building or outdoor hiking with large elevation changes.
Camera Image Embeddings.  Beyond the head pose trajectory derived from visual SLAM, the ego-centric camera images offer additional valuable information. For example, when a body part becomes intermittently visible, the image provides a strong cue of the wearer’s pose. However, direct utilization of the image content proves less useful, as it may capture distracting texture details when all we need is high-level semantics such as ”the left hand is above the waist.” Empirically, we found the CLIP embeddings [56], $E_{I}(\bm{I})$, provide a significant performance boost to the learning process while avoiding overfitting to superficial image characteristics.
@@
It is important to emphasize that embeddings provided by human-related backbones, such as networks trained for pose reconstruction from monocular videos, will not work well in our case. Figure 2 shows a typical input camera sequence when parts of the self-body (in this example, the lower arms) are visible. We can immediately notice the difference from the problem setting of downward-facing egocentric cameras, where in our case most body parts are not visible most of the time. This discrepancy causes existing network backbones for full-body motion to fail, and it may be tempting to assume that such image input might not be useful for full-body motion reconstruction. However, high-level descriptions of the images that contain scene information, such as ”hand reaching to the sink (which is typically at a standard height)” or ”a person kicking a football (implicitly indicating that the wearer might also soon interact with the ball),” are actually quite useful for spatial reasoning of the wearer’s end effectors. We hypothesize that this observation explains why CLIP embeddings are advantageous in our unique problem setting.
SLAM Point Cloud Embeddings.  Visual SLAM algorithms identify static feature points in the environment (e.g. corners and edges of furniture) and aggregate them over time to build 3D maps. These points offer crucial environment features to constrain motion generation, akin to pre-scanned scenes utilized in prior work [18, 38]. At each frame, we only consider the available SLAM feature points $\bm{S}$ within a 2m x 2m x 2m volume. The center of the volume is the current device position offset downwards by one meter, similar to prior works [60]. This ensures the model focuses only on relevant spatial information as the wearer moves around. To better handle the noisy and often incomplete nature of SLAM point clouds, we pre-train an autoencoder on the voxelized SLAM point clouds $V(\bm{S})$ within the bounding volume on all frames in our training dataset and use its encoder $E_{S}(\cdot)$ to generate point cloud embeddings $E_{S}(V(\bm{S}))$. While a new map may not offer much information right away, rich point cloud features could quickly build up if the wearer stays in the same environment for a prolonged period (e.g. 15 min) or if they have access to a prebuilt map.
@@ 3.2 Conditional Motion Diffusion Model
Online Inference of Long Sequences.  Our motion Diffusion model generates up to 4 seconds of motion ($T=240$ frames). To extend this for longer, coherent motions, previous research [7, 80, 58] suggests generating overlapping windows and enforcing consistency at overlaps during denoising. However, for online generation, we need to remove the dependency on future windows by using an autoregressive approach [23], where each window depends only on the previous one.
Figure 3: Autoregressive in-painting is performed at each reverse Diffusion step to allow long sequence generations both in high- and low-latency settings.
Specifically, when two windows overlap by $T-h$ frames (i.e., the current window advances by a stride of $h$), we enforce consistency during each of the $\bar{S}$ denoising steps. After each model evaluation $\hat{\bm{x}}_{0}=D(\bm{x}_{t},\bm{c},\bm{\tau}_{i})$, the prediction $\hat{\bm{x}}_{0}$ is overwritten by the overlapping prediction from the preceding window:
@@
where $\bm{m}\in\mathbb{R}^{T\times F}$ is a constant mask that is zero for the initial $T-h$ frames and one for the last $h$ frames. $\hat{\bm{x}}_{s0}=\textrm{cat}(\bm{x}^{-}_{0}[h\textrm{:}T],\bm{0}^{h\times F})$ denotes the prediction from the previous window, shifted by $h$ frames. $\odot$ denotes element-wise multiplication. Following this inpainting operation, we proceed to the denoising step with the updated $\hat{\bm{x}}_{0}$ using Eq. 2. We report the main results of our system with stride $h=180$.
However, eliminating the need for future windows is insufficient for online inference with minimal latency since a new window of motion is generated only every $h$ frames, resulting in a latency of $(h-1)\times\delta t$, where $1/\delta t$ is the frame rate. We additionally report our system’s results with $h=10$, indicating a latency of just $0.17$ seconds close to online requirements. Nonetheless, a smaller $h$ compromises motion quality, as it limits the use of future input conditions. In general, $h$ can be a tunable parameter to trade off quality and latency.
@@ 4 Experiments
Datasets and Experiment setup.  To address the limitation of evaluating on synthetic or smaller real-world datasets, we train and evaluate our system on a large-scale, first-of-its-kind real-device dataset Nymeria [45]. This dataset contains paired multi-modal HMD input signals (captured by Project Aria Glasses [14]) and ground-truth full-body motions (with Xsens inertial motion capture system [53]). The dataset covers a diverse range of daily activities and is around 300 hours in size. After initial filtering, we split the data into train, validation, and test split with 202, 3, and 56 hours of data correspondingly. We make sure all subjects and environment scenes in the test split are unseen during training, and distributions of subjects’ body sizes and activity scripts are roughly unbiased across the test split. We trained our models with a context window of 240 frames (4 sec) for 20 epochs or 3.5 days with 4 GPUs. The inference is done on a single Nvidia A100 GPU and achieves better than realtime throughput of $>70$ FPS with an online 0.17s-latency ($h=10$) model and $>1350$ FPS with high-latency (3s, $h=180$) model.

### v2

@@ Abstract
This paper investigates the generation of realistic full-body human motion using a single head-mounted device with an outward-facing color camera and the ability to perform visual SLAM. To address the ambiguity of this setup, we present HMD2, a novel system that balances motion reconstruction and generation. From a reconstruction standpoint, it aims to maximally utilize the camera streams to produce both analytical and learned features, including head motion, SLAM point cloud, and image embeddings. On the generative front, HMD2 employs a multi-modal conditional motion diffusion model with a Transformer backbone to maintain temporal coherence of generated motions, and utilizes autoregressive inpainting to facilitate online motion inference with minimal latency (0.17 seconds). We show that our system provides an effective and robust solution that scales to a diverse dataset of over 200 hours of motion in complex indoor and outdoor environments.
@@ 1 Introduction
Wearable devices such as smart glasses promise to become the cornerstone of next-generation personal computing. A key challenge is accurately interpreting the wearer’s motion from the device’s limited input signals, taking into account the social and environmental context at the moment. The capability to generate full-body movements solely from a single head-mounted device (HMD) in real-time, outdoors and indoors, will open the door to many downstream applications, including telepresence, fitness and health monitoring, and navigation.
@@
In this paper, we introduce HMD2 (Human Motion Diffusion from HMD), the first system, to our knowledge, capable of online generation of full-body movements from a single HMD (Project Aria Glasses [14]), conditioned on outward-facing egocentric camera streams in diverse environments. Given that such devices provide limited observation of the body and surroundings, the critical question is how to maximally utilize the input. Our approach reuses input data to generate features across different modalities, covering independent aspects of the environment and motion. Specifically, from the input streams, we mix and match analytical and learning toolboxes to extract 1. wearer’s head motion from off-the-shelf real-time visual SLAM; 2. environment feature points as a by-product of SLAM, important for motion disambiguation in complex scenes; and 3. head camera image embeddings (e.g. using CLIP [56]) for additional scene clues and intermittently visible body parts.
However, full recovery of the wearer’s motion is still highly under-constrained, given our input. Our system adopts a generative approach with a diffusion-based Transformer backbone to balance motion reconstruction and generation, enabling diverse outcomes, such as varying leg movements, from the same inputs. Additionally, our diffusion model can predict motions with minimal future information (0.17 s), supporting online and real-time use cases.
Contrary to evaluations using large synthetic datasets or small-scale real-world datasets, we train and test our system on the extensive 200-hour real-world Nymeria dataset [45] recorded with publicly available head-mounted device, containing various indoor and outdoor activities performed by over 100 subjects with diverse body sizes and demographics. While most existing research on motion tracking is evaluated solely based on reconstruction accuracy, we acknowledge the inherent ambiguity in our problem and evaluate our system on generation fidelity and diversity as well. Our contributions are summarized as follows:
@@
    We demonstrate the adaption of a time-series motion diffusion model for online autoregressive inference through inpainting, eliminating the dependency on future sensor input and achieving minimal latency.
@@ 2 Related Work
Human Motion from Sparse Sensors.  Capturing motion with wearable sensors has gained interest across fields like Computer Vision, Graphics, and Health. Self-contained sensors like IMUs [66], electromagnetic sensors [35], and EMGs [10] offer motion reconstruction without the need for costly studios with multiple cameras. The sparse sensor placement reduces user friction, but high noise levels require learning methods to improve reconstruction. Examples include six IMUs configurations [66, 28, 78, 33, 77], head and wrists VR trackers [32, 13, 90, 7, 74, 31], and hybrid approaches with an external RGB camera [75].
Our approach uses a single wearable device to minimize user friction, though this complicates the recovering of motion. However, for many applications like telepresence, visually appealing, realistic, and diverse inferred motions are often more important than precision. Thus, we evaluate our system not just on reconstruction accuracy but also on realism and diversity – metrics often overlooked in this field.
Pose and Motion from Egocentric Cameras.  Wearable egocentric cameras are ideal for self-contained motion generation systems, which saw increasing research interest. Two main types of body-mounted cameras – downward-facing (often fish-eye) and outward-facing – have been the focus of research. Most studies on downward cameras [57, 8, 61, 68, 73, 85, 87, 51, 34, 41], directly predict current pose from corresponding images, sacrificing temporal coherence. Wang et al. [67] addressed this by adopting a diffusion model for temporal regularization in a separate refinement stage, which inspired us to adopt a diffusion backbone and a single-stage time-window-based learning architecture. Both synthesized [4] and, recently, real-device [3] datasets are used to train and evaluate such methods.
Outward-facing cameras are more common on current devices (e.g. Project Aria [14]), though egocentric motion generation is less explored in this setup. A key challenge identified in early work with chest-mounted cameras [30] is intermittent body visibility, which makes the task underconstrained. Later works [80, 81, 43] explored simulation methods that leveraged physics to address missing motion information. EgoEgo [39] demonstrated the generalizability of single camera systems to large-scale datasets. We build upon EgoEgo, while utilizing additional visual cues beyond head pose inference and enhancing support for non-flat terrains and low-latency long sequence generation.
@@
Learning-based Pose and Motion Generation.  Generating controllable and realistic human movements is a long-standing goal in computer graphics and vision. Modern deep learning opens new possibilities for this problem, with earlier attempts exploring both regression-based [25, 24] and generative [21, 40] frameworks. Recently diffusion models demonstrated impressive capabilities in the generative setting across various tasks such as text [58, 91], music [62], and audio [5] conditioned motion generation. While the field starts to see conditional diffusion methods where the control signal is temporally dense [7, 13], frameworks that generate motions in an online fashion with minimal latency [64] are still underexplored. Our work adopts autoregressive inpainting for low-latency inference – this concept of autoregressive diffusion models has been explored in the motion domain albeit in different contexts [19, 86, 59, 79].
The success of diffusion models in motion synthesis has also intrigued researchers to use them for pose reconstruction, e.g. from third-person view, especially when ambiguity exists [11, 12, 15, 16, 26, 84, 83]. Our task is highly ambiguous as well, and our system adopts Transformer-based diffusion models to generate temporally coherent motions.
@@ 3 Method
Figure 1: Overview: HMD2 generates realistic full-body motion that aligns with the signals from a single head-mounted device. Using the image streams from the egocentric camera and head trajectory with the feature cloud from the onboard SLAM system, we employ a diffusion-based framework to generate the wearer’s full-body motion.
@@
Given the under-constrained nature of the task, we employ a diffusion model [22] with a time-series Transformer encoder [65] to model the motion distribution. To ensure temporal consistency during streaming, we use autoregressive inpainting during denoising, aligning new body motion with previous predictions.
@@ 3.1 Multi-modal Scene and Motion Conditions
Head Pose Trajectory.  The device pose provides precise spatial location and movement of the wearer’s head. We augment the device pose vector with its linear and angular velocity vector $(\bm{v},\bm{\omega})$ computed from finite differences to form $\bm{p}=\{\bm{t},\bm{R},\bm{v},\bm{\omega}\}$. We canonicalize each window of $\{\bm{p}\}_{0,1,\cdots,T}$ to its first frame $\bm{p}_{0}$, allowing the model to function in arbitrarily large spaces and generate infinitely long sequences. This is crucial for navigation in a multistory building or outdoor hiking with large elevation changes.
Camera Image Embeddings.  Beyond the head pose trajectory derived from SLAM, the egocentric camera images offer additional valuable information. For example, when a body part becomes visible, the image provides a strong cue of the wearer’s pose. However, direct utilization of the image content proves less useful, as it may capture distracting texture details when all we need is high-level semantics such as ”the left hand is above the waist.” Empirically, we found that CLIP embeddings [56], $E_{I}(\bm{I})$, provide significant performance boost to the learning process while avoiding overfitting to superficial image characteristics.
@@
It is crucial to note that embeddings from human-related backbones, such as those trained for pose reconstruction from monocular videos, do not perform well in our case. Figure 2 shows a typical input camera sequence when only a few parts of the body (hands in this case) are visible. This differs significantly from downward-facing egocentric cameras, which observe most of the body. This discrepancy leads to failures in existing network backbones for full-body motion, and it may be tempting to assume that such input might not be useful for full-body motion reconstruction. However, high-level descriptions of the images that contain scene information, such as “hand reaching to the sink (which is typically at a standard height)” or “a person kicking a football (implicitly indicating that the wearer might also soon interact with the ball)”, are actually quite useful for spatial reasoning of the wearer’s end effectors. We hypothesize that this observation explains why CLIP embeddings are advantageous in our unique problem setting.
SLAM Point Cloud Embeddings.  Visual SLAM algorithms identify static feature points in the environment (e.g. corners and edges of furniture) and aggregate them over time to build 3D maps. These points offer crucial environment features to constrain motion generation, akin to pre-scanned scenes utilized in prior work [18, 38]. At each frame, we only consider the SLAM feature points $\bm{S}$ within a 2m x 2m x 2m volume. The center of the volume is the current device position offset downwards by one meter, similar to prior works [60]. This ensures the model focuses only on relevant spatial information as the wearer moves around. To better handle the noisy and often incomplete SLAM point clouds, we pre-train an autoencoder on the voxelized SLAM point clouds $V(\bm{S})$ within the bounding volume on all frames in our training dataset and use its encoder $E_{S}(\cdot)$ to generate point cloud embeddings $E_{S}(V(\bm{S}))$. While a new map may not offer much information right away, rich point cloud features could quickly build up if the wearer stays in the same environment for a prolonged period (e.g. 15 min) or if they have access to a prebuilt map.
@@ 3.2 Conditional Motion Diffusion Model
Figure 3: Autoregressive inpainting is performed at each reverse diffusion step to allow long sequence generations both in high- and low-latency settings.
Online Inference of Long Sequences.  Our motion diffusion model generates up to 4 seconds of motion ($T=240$ frames). To extend this for longer, coherent motions, previous research [7, 82, 58] suggests generating overlapping windows and enforcing consistency at overlaps during denoising. However, for online generation, we need to remove the dependency on future windows by using an autoregressive approach [23], where each window depends only on the previous one. Specifically, when two windows overlap by $T-h$ frames (i.e., the current window advances by a stride of $h$), we enforce consistency during each of the $\bar{S}$ denoising steps. After each model evaluation $\hat{\bm{x}}_{0}=D(\bm{x}_{t},\bm{c},\bm{\tau}_{i})$, the prediction $\hat{\bm{x}}_{0}$ is overwritten by the overlapping prediction from the preceding window:
@@
where $\bm{m}\in\mathbb{R}^{T\times F}$ is a constant mask that is zero for the initial $T-h$ frames and one for the last $h$ frames. $\hat{\bm{x}}_{s0}=\textrm{cat}(\bm{x}^{-}_{0}[h\textrm{:}T],\bm{0}^{h\times F})$ denotes the prediction from the previous window, shifted by $h$ frames. $\odot$ denotes element-wise multiplication. Following this inpainting operation, we move to denoising step with the updated $\hat{\bm{x}}_{0}$ using Eq. 2. We report the main results of our system with stride $h=180$.
However, eliminating the need for future windows is insufficient for online inference with minimal latency since a new window of motion is generated only every $h$ frames, resulting in a latency of $(h-1)\times\delta t$, where $1/\delta t$ is the frame rate. We additionally report our results with $h=10$, indicating a latency of just $0.17$ seconds, close to online requirements. Nonetheless, a smaller $h$ compromises motion quality, as it limits the use of future information. In general, $h$ can be a tunable parameter to trade off quality and latency.
@@ 4 Experiments
Datasets and Experiment setup.  To address the limitation of evaluating on synthetic or smaller real-world datasets, we train and evaluate our system on a large-scale, first-of-its-kind real-device dataset Nymeria [45]. This dataset contains paired multi-modal HMD input signals (captured by Project Aria Glasses [14]) and ground-truth full-body motions (with Xsens inertial motion capture system [52]). The dataset covers a diverse range of daily activities and is around 300 hours in size. After initial filtering, we split the data into train, validation, and test split with 202, 3, and 56 hours of data correspondingly. We make sure all subjects and environment scenes in the test split are unseen during training, and distributions of subjects’ body sizes and activity scripts are roughly unbiased across the test split. We trained our models with a context window of 240 frames (4 sec) for 20 epochs or 3.5 days with 4 GPUs. The inference is done on a single Nvidia A100 GPU and achieves better than real-time throughput of $>70$ FPS with an online 0.17s-latency ($h=10$) model and $>1350$ FPS with high-latency (3s, $h=180$) model.